# INQUIRE-Search: A Framework for Interactive Discovery in Large-Scale Biodiversity Databases

### v2

@@ 2.2.2 Methodology
    For each image, a researcher marked informative images containing the forest community of interest (i.e., an individual of the target forest type could be clearly distinguished). We allowed individual images to be marked as informative in multiple searches if the image contained individuals representing both forest types (e.g. a coniferous and deciduous tree seedling). While processing images for searches related to trees, we also decided to include trees and shrubs within the same target forest type as it was difficult to distinguish young trees from shrubs in images.
  5. 5.
    We extracted MTBS burn severity categories for the selected image locations and related burn severity to regeneration outcomes. Burn severity categories were aggregated from the native 30-m Landsat resolution to 0.01 degree resolution (Figure˜5) or consistency with the exported image coordinates. We aggregated burn severity categories using the most common burn severity category among contributing pixels using the terra package in R4 [hijmans2022package].
  6. 6.
    We count the number of marked images for each forest type and burn severity category.
2.2.3 Results and Analysis
Figure 6: (Top) Query & Verification. Top INQUIRE-Search returns for "Young coniferous trees in burned forest" with incorrect or irrelevant images removed through expert review. The images were marked correct if the right type of tree is visible in any part of the image, even if not in the foreground. (Bottom) Tree observations across burn severity. Distribution of observed young coniferous and deciduous trees in community science collected images show strong relationships with burn severity regions.
Of the 200 images exported from INQUIRE-Search for each search, there were 100 (coniferous) and 112 (deciduous) images with reported coordinates within the High Park burn perimeter. We marked more informative images for deciduous forest species (78) than coniferous (45). We also marked five images depicting both coniferous and deciduous individuals, indicating areas that could recover into mixed forest communities
Among images marked for analysis, we found strong relationships between burn severity and post-fire regeneration of both coniferous and deciduous forest types. For both forest types, most marked images of regenerating vegetation occurred in areas classified as unburned-to-low or low burn severity, with few observations in moderate or high severity areas (Figure˜6, left). However, coniferous regeneration showed a stronger negative association with increasing burn severity: 93.3% of coniferous images were located in unburned-to-low or low severity areas, with only 4.4% and 2.2% in moderate and high burn severity areas, respectively (Figure˜6, right). In contrast, deciduous regeneration was somewhat more evenly distributed, with 84.6% of marked images in unburned-to-low or low severity areas, and a higher proportion (7.6% each) in areas burned at moderate and high severity (Figure˜6, right). These patterns suggest that post-fire forest regeneration—or at least regeneration captured by citizen images—is more common in lower severity areas, with coniferous species showing greater sensitivity than deciduous species.
2.2.4 Findings
In this case study, we demonstrate that INQUIRE-Search can be used to effectively identify post-fire forest regeneration patterns across burn severity gradients. Our approach successfully distinguished between regenerating coniferous and deciduous forest types and provided spatially explicit insights into post-disturbance recovery within the first decade after the High Park fire. We found that regeneration was most commonly observed in unburned areas or areas burned at low severity, with coniferous species showing greater sensitivity to increasing burn severity than deciduous species. This is consistent with previous research documenting that increased fire severity decreases conifer regeneration likelihood both regionally and specifically after the High Park fire [davis2023reduced, wright2017post]. Overall, this method offers a scalable and accessible tool for monitoring ecological recovery using public image databases.
Despite INQUIRE-Search’s utility for harnessing citizen science data and identifying post-fire vegetation communities, several limitations remain. Image coordinates are rounded to 0.01° to match the limiting spatial precision of many iNaturalist records, but this prevents fine-scale matching to burn severity or site attributes such as seed source or soil moisture. Classification from user-submitted photographs also introduces uncertainty: landscape images often obscure species identification, and deciduous seedlings, saplings, shrubs, and herbaceous plants can appear morphologically similar. Our expert reviewer frequently could not distinguish these groups, reflecting a broader challenge shared with remote-sensing approaches: early-seral vegetation often remains visually and spectrally ambiguous for years after fire. Finally, iNaturalist data are opportunistic and biased toward accessible, aesthetically appealing areas, which may result in severely burned sites being underrepresented (Fig. 3). Thus, while INQUIRE is a powerful supplemental tool for assessing post-fire recovery, its results should be interpreted alongside field data and remote-sensing derived products to minimize bias.
2.3 Wildlife Mortality
@@
Figure 7: (Left) Wildlife mortality trends vary by season and location. Collisions with human-made structures such as windows are a major cause of bird deaths, varying seasonally and with urbanization. (Right) Comparing avian mortality in rural vs urban regions. Using INQUIRE-Search, we compared image-based evidence of avian mortality between urban (Boston, MA) and rural (Pioneer Valley, MA) sites to examine how seasonal risks differ across anthropogenic contexts.
2.3.1 Task Description
Understanding patterns of avian mortality is fundamental to conservation efforts, particularly given the dramatic decline of nearly 3 billion birds in North America since 1970 [rosenberg2019decline]. Mortality risk varies seasonally as birds encounter different threats across the annual cycle [loss2015direct, marra2015call], and these patterns differ across landscapes due to variation in local hazards and bird abundance. Collisions with buildings—one of the leading human-driven causes of bird death [loss2014estimation]—are strongly seasonal, peaking during migration [riding2021multi, scott2023causes], and are further shaped by anthropogenic context, with tall structures in rural areas often generating disproportionately high collision rates [hager2017continent]. Understanding how mortality patterns shift across space and time is therefore essential for designing targeted interventions and addressing ongoing population declines.
However, documenting spatio-temporal patterns of animal mortality presents significant methodological challenges. Direct observation of dead animals in the wild is relatively rare because mortality events occur infrequently and alternative approaches such as marking or tracking individuals are labor-intensive, expensive, and only feasible for certain taxa at limited scales [yanco2025tracking]. Leveraging large-scale image datasets to identify where and when animals die would fill key gaps in our ability to study the patterns and processes of wildlife mortality.
In this case study, we used INQUIRE-Search to investigate seasonal avian mortality dynamics in an urban (Boston, MA) and rural (Pioneer Valley, near Amherst, MA) setting and compare relative mortality dynamics between the two sites. We hypothesized that (1) mortality rates would be higher in urban contexts, due to the greater density of man-made infrastructure that poses a collision risk [loss2015direct] and (2) mortality rates would be greatest during seasonal migrations when avian densities are highest [scott2023causes].
2.3.2 Methodology
We used a systematic approach to estimate mortality rates at both sites, by querying INQUIRE-Search:
  1. 1.
    We queried INQUIRE-Search for "dead bird" in each of the two equal-sized bounding boxes. We visually inspected results and removed records of live birds, single feathers, non-avian species, and duplicate observations of the same mortality event (duplicates defined by shared species, month, latitude, and longitude).
  2. 2.
    To compare the number of observed mortalities to the overall number of iNaturalist observations, we also downloaded the total number of iNaturalist observations of birds within those spatial bounding boxes using the rinat package in R [barve2022rinat].
  3. 3.
    We summarized both mortalities and observations by month for each study area and then calculated monthly mortality "rate" as the monthly mortality count divided by monthly observation count. Because we expect the baseline bird abundance, observation effort, and mortality rates to differ among sites, we normalized each month’s effort by dividing it by the mean monthly site-specific mortality rate and then taking the $\log_{2}$. This results in a "Mortality Index" which controls for differences in observer effort and underlying bird population abundances. Taking the log of mortality rate symmetrically centers the index on zero such that values of this index above 0 represent the fold-change above mean mortality rates whereas values below 0 reflect the converse.
@@ 2.3.4 Findings
However, several important limitations must be considered when interpreting these results. First, detection biases likely vary substantially between urban and rural environments. In urban settings, dead birds are more likely to be discovered on sidewalks, building perimeters, and other high-visibility locations where citizen scientists frequently travel. On the other hand, rural mortality events may be significantly underdetected due to the tendency for carcasses to occur in forested or remote areas where they are quickly removed by scavengers or decompose unnoticed. Second, the observer demographics and motivations for using iNaturalist may differ between urban and rural areas, potentially introducing additional sampling biases. Third, our mortality index assumes that the ratio of dead bird observations to total bird observations provides a meaningful proxy for actual mortality rates, but this assumption may not hold if detection probabilities vary systematically across seasons, locations, or species.
2.4 Resolving Plant Phenology
2.4.1 Task Description
Figure 9: Four pheonological stages of common milkweed. Example observations identified using INQUIRE-Search that correspond to the four phenological stages analyzed in this study: emergence, flowering, seeding, and senescence.
Studying how the timing of specific life history events is changing is a primary way researchers are detecting biodiversity response to ongoing global change [parmesan2003globally]. Often highly malleable, phenological shifts in the timing of events like plant germination and reproduction, can serve as early warning signs that precede other biodiversity responses like demographic, genetic, or geographic shifts which can threaten the persistence of communities.
Traditionally, assessing phenological changes involves a significant amount of field study, to record and track the timing of specific events year after year. For example, Austin et al. (2024) [austin2024climate] use multiple phenology datasets gathered over 5 years comprising weekly transects during growing months (March-November) to understand shifts in flowering onset and duration in a single plot in the Ozark Highlands in Missouri, USA. While these sorts of datasets permit fine scale assessments of how species in a specific community are responding to climate fluctuations, detecting and attributing phenological change at broader spatial scales, which can enable a more holistic view of biodiversity change [doi2017macroecological], remains challenging due to widespread data limitations. Some macrophenological studies have managed to overcome these data limitations by using computer vision approaches to extract phenological information from a variety of publicly available data (e.g., digitized herbarium specimens) [williamson2025long]. Yet, most macrophenological studies focus on relatively simple metrics to quantify phenological shifts, such as mean flowering date and flowering duration, as limited intra-annual scale data often precludes more detailed assessments of shift in phenological stages like emergence, flowering onset, seeding, and senescence.
Now however, the wealth of photo-based biodiversity observations produced via community science platforms like iNaturalist can be leveraged to overcome these data shortfalls, if we can extract phenological information from observations in a systematic and reliable manner. Here we test whether INQUIRE-Search can be leveraged to piece together the full life history of an ecologically significant species, common milkweed (Asclepias syriaca), in a specific region (Southern Quebec). We show that INQUIRE-Search can successfully identify observations that cluster into specific phenological stages, emergence, flowering, seeding, and senescence (Figure˜9).
2.4.2 Methodology
We used INQUIRE-Search to retrieve images of target species across multiple life stages:
  1. 1.
    We queried INQUIRE-Search with specific queries for each phenological stage: (1) emergence (“Milkweed germinating or emerging”), (2) flowering (“Milkweed flowering”), (3) seeding (“Milkweed producing seeds or milkweed with seeds”), and (4) senescence (“Milkweed dying or withering or senescence”).
  2. 2.
    We used both a species filter (“Asclepias syriaca”) and a geographic filter (Latitude: 45.03 to 46.54 degrees, Longitude: -74.68 to -71.66 degrees) to isolate observations of common milkweed in southern Quebec.
  3. 3.
    We then used a set of criteria to identify observations that corresponded to specific phenological stages. For emergence, we considered observations with milkweed plants containing less than four pairs of adult leaves. For flowering, we considered observations with open petals. For seeding, we considered observations with open seed pods and visible seeds. For senescence, we considered observations with milkweed plants that no longer exhibited green leaves. For each stage, we queried 200 observations using INQUIRE-Search and marked the observations which corresponded to the matching phenological stage.
@@ 2.4.2 Methodology
Figure 10: (Top) Query & Verification. Top INQUIRE-Search returns for "flowering" phenological stage with incorrect or irrelevant images removed through expert review. (Bottom) Phenological stages occur at significant times throughout the year. (left) density of observations from INQUIRE-Search corresponding to different stages. (right) distribution of observations from INQUIRE-Search in different phenological stages visualized with results from ANOVA/Tukey test.
2.4.3 Results and Analysis
@@ 2.4.4 Findings
This case study shows that INQUIRE-Search can be used systematically to resolve detailed phenological information at regional scales for specific species. Some phenological stages (flowering and seeding) were more successfully resolved than others (emergence and senescence). While this could be a limitation of the models ability to identify relevant observations to those stages (the difference between an emerging plant and a young plant is minor) it is also likely due to bias in the community science observations queried, as observations of flowers and seeds are likely more common as they can aid taxonomic identification [iNaturalist202xHighQualityObservations]. Resolving patterns of senescence will also require specific considerations to account for the fact that senescent plants can also be observed long after senescence, evidenced by the fact we identified observations of senescent milkweed throughout the year. Nonetheless, our findings point to the potential utility of INQUIRE-Search to benefit our ability to resolve detailed phenological stages. Considering this sort of fine0scale phenological information would have traditionally required daily or weekly surveys to resolve at the level of a single site, let alone a region like Southern Quebec, our results demonstrate that INQUIRE-Search has the potential to drastically increase our ability to extract phenological information from the wealth of rapidly growing community science observations.
2.5 Whale Re-Identification
2.5.1 Task Description
Animal re-identification (re-ID)—the ability to recognize individual animals across space and time—is central to estimating population size and survival, reconstructing migratory routes, evaluating site fidelity, and monitoring long-term shifts in habitat use [krebs1989ecological]. For wide-ranging marine mammals such as humpback whales, these insights depend on assembling longitudinal photo-ID datasets that link repeated sightings of the same individual across years and ocean locations [katona1981identifying, calambokidis2001movements, martin1984migration]. The pigmentation and trailing-edge patterns on the ventral surface of the tail fluke are often used for identification, and large curated datasets like HappyWhale [howard2018humpback] have been built around these features.
However, scaling re-ID efforts require two components: (1) locating suitable images that clearly show the required body region, and (2) matching those images to known individuals using automated or semi-automated algorithms. While considerable progress has been made on the latter components, with deep re-ID models having shown strong potential for automating fluke matching, locating identifiable images remains a major bottleneck [otarashvili2024multispecies]. Although many fluke images exist in open platforms such as iNaturalist, they are seldom used because there is no effective way to retrieve only those showing a clear, properly oriented ventral fluke suitable for re-identification, while filtering out low-quality or improperly positioned images (see Figure˜11). These identifiable fluke images are scattered across various sources, from community science platforms to tourism photography, each with varying quality, pose, lighting, and availability of metadata.
INQUIRE-Search addresses this gap by enabling targeted retrieval of individually-identifiable images from large, unstructured image collections. Using queries such as "white underside of humpback whale fluke", species filters, and relevancy ranking, the system can surface high-quality fluke images from a global pool of observations. This creates the possibility of augmenting curated catalogs with previously inaccessible sightings, expanding geographic and temporal coverage, and supporting dynamic population monitoring across ocean basins. In this case study, we assess whether INQUIRE-Search can (1) retrieve identifiable humpback whale fluke images from iNaturalist and (2) match them to individuals in the HappyWhale dataset.
2.5.2 Methodology

### v3

@@ 2.3.2 CS2: Post-fire Forest Regrowth
Verification: An image was marked informative when it contained the forest community of interest (i.e., an individual of the target forest type could be clearly distinguished). We allowed individual images to be marked as informative in multiple searches if the image contained individuals representing both forest types (e.g. a coniferous and deciduous tree seedling). While processing images for searches related to trees, we also decided to include trees and shrubs within the same target forest type as it was difficult to distinguish young trees from shrubs in images. Top 200 images were inspected.
Analysis: Coordinates of informative images were combined with MTBS burn severity categories by aggregating the 30 m burn data to the 0.01 precision typical of iNaturalist coordinates and assigning each location the modal severity class using the terra package in R4 [hijmans2022package]. We then counted marked images by forest type and burn severity to evaluate how regeneration outcomes varied across burn categories.
2.3.3 CS3: Wildlife Mortality
@@
Figure 4: (Left) Wildlife mortality trends vary by season and location. Globally, collisions with human-made structures such as windows are a major cause of bird deaths. (Right) Comparing avian mortality in rural vs urban regions. Using INQUIRE-Search, we compared image-based evidence of avian mortality between urban (Boston, MA) and rural (Pioneer Valley, MA) sites to examine how seasonal risks differ across anthropogenic contexts.
Overview: Understanding patterns of avian mortality is fundamental to conservation efforts, particularly given the loss of nearly three billion birds in North America since 1970 [rosenberg2019decline]. Mortality risk varies seasonally as birds encounter different threats across the annual cycle [loss2015direct, marra2015call], and collisions with human-made structures–one of the leading anthropogenic causes of bird death [loss2014estimation]–show strong seasonal peaks during migration [riding2021multi, scott2023causes] and vary across landscapes due to differing local hazards and bird abundances [hager2017continent]. Yet documenting spatiotemporal patterns of mortality is challenging because mortality events are rare, and alternative monitoring methods such as marking or tracking individuals are resource-intensive and limited in scale [yanco2025tracking]. Since large-scale image repositories could reveal where and when birds die, this case study uses INQUIRE-Search to investigate seasonal mortality dynamics in an urban (Boston, MA) and rural (Pioneer Valley, MA) setting and compare relative mortality between these contrasting anthropogenic contexts (Fig 4).
Search Strategy: We queried for "dead bird" filtering by each of the two equal-sized bounding boxes using location filters. We use latitude (42.31 to 42.38 N) and longitude (-71.14 to -71.01 W) for the urban study area, and latitude (42.31 to 42.38 N) and longitude (-72.64 to -72.51 W) for the rural study area.
Verification: We excluded images of live birds, single features, non-avian taxa, and duplicate observations of the same event (duplicate sharing species, month, latitude and longitude) to identify true avian mortality events.
Analysis: To control for differing bird abundance, observation effort, and mortality among sites, we calculate the mortality index defined as:
$\text{MortalityIndex}_{m,s}=\log_{2}\!\left(\frac{R_{m,s}}{\overline{R}_{s}}\right),$
where the monthly mortality rate is
$R_{m,s}=\frac{D_{m,s}}{O_{m,s}},$
and the mean monthly mortality rate for site $s$ is
$\overline{R}_{s}=\frac{1}{12}\sum_{m=1}^{12}R_{m,s}.$
Here, $D_{m,s}$ is the mortality count and $O_{m,s}$ is the observation count for month $m$ at site $s$. Values above zero indicate fold increases in mortality relative to the site’s annual mean, while values below zero indicate fold decreases. The total observation count was extracted using the rinat package in R [barve2022rinat].
2.3.4 CS4: Resolving Plant Phenology
Overview: Studying the timing of life-history events is a primary way researchers detect biodiversity responses to global change [parmesan2003globally], as phenological shifts in processes such as germination, flowering, and senescence can serve as early warning signals that precede demographic, genetic, or geographic change. While fine-scale phenological datasets collected through repeated field surveys provide valuable local insight [austin2024climate], detecting and attributing phenological change at broader spatial scales remains challenging due to limited temporal resolution and data availability [doi2017macroecological]. Although some macrophenological studies have used computer vision to extract phenological information from digitized specimens [williamson2025long], most rely on coarse metrics such as mean flowering date because intra-annual data on distinct phenophases are sparse. The growing volume of geotagged plant observations on community-science platforms like iNaturalist offers a path to overcoming these limitations, provided phenological stages can be systematically identified from photographs. In this case study, we test whether INQUIRE-Search can recover detailed phenophases—emergence, flowering, seeding, and senescence—for common milkweed (Asclepias syriaca) in southern Québec (Fig 5).
Figure 5: Four phenological stages of common milkweed. Example observations identified using INQUIRE-Search that correspond to the four phenological stages analyzed in this study: emergence, flowering, seeding, and senescence.
Search Strategy: We created stage-specific text queries targeting (1) emergence ("Milkweed germinating or emerging"), (2) flowering ("Milkweed flowering"), (3) seeding ("Milkweed producing seeds or milkweed with seeds"), and (4) senescence ("Milkweed dying or withering or senescence"). We also used a species filter ("Asclepias syriaca") and a geographic filters for latitude (45.03 to 46.54 N) and longitude (-74.68 to -71.66 W) to limit observations to southern Quebec.
Verification: Top 200 images were labeled as informative according to strict morphological criteria:
  • Emergence: <4<4 pairs of adult leaves
  • Flowering: presence of open petals
  • Seeding: open pods with visible seeds
  • Senescence: absence of green leaves
Analysis: Observation dates were converted to Day-of-Year (DOY). Mean DOY values were compared among stages using ANOVA, followed by Tukey’s HSD to identify pairwise differences.
2.3.5 CS5: Whale Re-Identification
Overview: Animal re-identification – the ability to recognize individual animals across space and time – is essential for estimating population size and survival, reconstructing migratory routes, evaluating site fidelity, and monitoring long-term shifts in habitat use across different species and ecosystems [krebs1989ecological]. For wide-ranging marine mammals such as humpback whales, these insights depend on longitudinal photo-ID datasets linking repeated sightings across years and ocean basins [katona1981identifying, calambokidis2001movements, martin1984migration], typically using pigmentation and trailing-edge patterns on the ventral fluke [howard2018humpback]. Although many fluke images exist in open repositories like iNaturalist, scaling re-ID remains constrained by the difficulty of locating images that show a clear, properly oriented fluke suitable for identification, as existing filters cannot reliably distinguish low-quality, mis-angled, or incomplete views (Fig 6) [otarashvili2024multispecies]. Because identifiable fluke photographs are scattered across diverse image sources with variable pose, lighting, and metadata, this case study evaluates whether INQUIRE-Search can surface individually identifiable humpback whale flukes from unstructured community-science imagery and match them to known individuals in the HappyWhale dataset.
Search Strategy: We queried for "white underside of humpback whale fluke" with the humpback whale species filter.
Verification: Images were considered informative only if the ventral fluke was clearly visible, unobstructed, and in focus. Top 200 retrievals were inspected.
Analysis: For analysis, we assess whether any iNaturalist retrievals corresponded to known individuals in the HappyWhale dataset. To accelerate matching, we cropped each image using Grounding DINO [liu2024grounding] and embedded both the iNaturalist and HappyWhale datasets using a multi-species re-identification model [otarashvili2024multispecies]. For each of the verified INQUIRE-Search images, we retrieved the top three closest HappyWhale candidates by embedding similarity and manually reviewed these candidates to confirm matches. Although not conducted by experts, humpback flukes possess distinct and easily recognizable pigmentation and trailing-edge patterns that allow reliable identification by non-experts.
3 Results
3.1 CS1: Seasonal Variation in Bird Diets
Results and Analysis
Our analysis revealed substantial variation in the availability of dietary information across the five study species shown in Fig 7. American Robin yielded the highest number of informative (i.e., containing dietary information) images (669 total), followed by Red-bellied Woodpecker and American Tree Sparrow (395 and 188, respectively), while for both Ancient Murrelet and Gray-cheeked Thrush we found a negligible number of images containing dietary information. Among the species for which dietary information was available, seasonal patterns in image availability were also evident: American Robin had more informative images in summer than winter, while Red-bellied Woodpecker and American Tree Sparrow had more images available for winter months.
Figure 7: (Top) Query & Verification. Filtering top INQUIRE-Search outputs for "Red-bellied Woodpecker" and "Invertebrate." (Bottom) Diet comparisons. INQUIRE-Search yields dietary patterns that align closely with those documented in the SAviTraits reference dataset.
For the three species with sufficient data, dietary composition derived from INQUIRE-Search closely matched reference data from SAviTraits (Fig 7). For example, in summer, American Robin exhibited an even split between fruit and invertebrate consumption (50% each) in both datasets, while Red-bellied Woodpecker showed a diet dominated by seeds (50% in INQUIRE-Search vs. 40% in SAviTraits), with smaller contributions from invertebrates, fruits, and plant material. In winter, all three species exhibited diets strongly dominated by a single food type, and we again observed close agreement between INQUIRE-Search outputs and SAviTraits data.
Findings
We demonstrate that our method can effectively reproduce dietary patterns documented in established literature, provided that sufficient image data are available. The close alignment between INQUIRE-Search-derived avian diets and SAviTraits reference data highlights the potential of community science imagery to yield accurate ecological insights when properly analyzed through a systematic search methodology. However, this approach is most effective for commonly photographed species such as American Robin and Red-bellied Woodpecker. Two of the species we examined lacked sufficient image data for meaningful analysis despite being well-studied in the scientific literature and represented in databases like SAviTraits, underscoring a fundamental constraint of community-sourced data. This limitation is likely even more pronounced for understudied species—precisely those for which new dietary data are most needed—because they are also underrepresented in iNaturalist and thus rarely retrievable with INQUIRE-Search.
3.2 CS2: Post-fire Forest Regrowth
Figure 8: (Top) Query & Verification. Filtering top INQUIRE-Search returns for "young coniferous trees in burned forest." (Bottom) Tree observations across burn severity. Distribution of observed young coniferous and deciduous trees in community science collected images show strong relationships with burn severity regions.
Results and Analysis
Of the 200 images exported from INQUIRE-Search for each search, there were 100 (coniferous) and 112 (deciduous) images with reported coordinates within the High Park burn perimeter. We marked more informative images for deciduous forest species than coniferous (78 vs. 45 images, respectively). We also marked five images depicting both coniferous and deciduous individuals, indicating areas that could recover into mixed forest communities.
Among images marked for analysis, we found strong relationships between burn severity and post-fire regeneration of both coniferous and deciduous forest types. For both forest types, most marked images of regenerating vegetation occurred in areas classified as unburned-to-low or low burn severity, with few observations in moderate or high severity areas (Fig 8, left). However, coniferous regeneration showed a stronger negative association with increasing burn severity: 93.3% of coniferous images were located in unburned-to-low or low severity areas, with only 4.4% and 2.2% in moderate and high burn severity areas, respectively (Fig 8, right). In contrast, deciduous regeneration was somewhat more evenly distributed, with 84.6% of marked images in unburned-to-low or low severity areas, and a higher proportion (7.6% each) in areas burned at moderate and high severity (Fig 8, right). These patterns suggest that post-fire forest regeneration, as captured by community-science images, is more common in lower severity areas, with coniferous species showing greater sensitivity than deciduous species.
Findings
In this case study, we demonstrate that INQUIRE-Search can be used to effectively identify post-fire forest regeneration patterns across burn severity gradients. Our approach successfully distinguished between regenerating coniferous and deciduous forest types and provided spatially explicit insights into post-disturbance recovery within the first decade after the High Park fire. Our findings are consistent with previous research documenting that increased fire severity decreases conifer regeneration likelihood both regionally and specifically after the High Park fire [davis2023reduced, wright2017post].
Several limitations remain. Image coordinates are rounded to 0.01 to match the limiting spatial precision of many iNaturalist records, but this prevents fine-scale matching to burn severity or site attributes such as seed source or soil moisture. Classification from user-submitted photographs also introduces uncertainty, as landscape images often obscure species identification, and deciduous seedlings, saplings, shrubs, and herbaceous plants can appear morphologically similar. Finally, iNaturalist data are opportunistic and biased toward accessible, aesthetically appealing areas, which may result in severely burned sites being underrepresented (Fig 3). Thus, while INQUIRE is a powerful supplemental tool for assessing post-fire recovery, its results should be interpreted alongside field data and remote-sensing derived products to minimize bias.
3.3 CS3: Wildlife Mortality
@@
However, several important limitations must be considered when interpreting these results. First, detection biases differ substantially across landscapes. Urban carcasses occur in high-visibility areas and are readily detected, while rural mortality events are often missed because they occur in remote or forested habitats where scavenging and decomposition quickly eliminate evidence. Additionally, our mortality index assumes that the ratio of dead bird observations to total bird observations provides a meaningful proxy for actual mortality rates, but this assumption may not hold if detection probabilities vary systematically across seasons, locations, or species.
3.4 CS4: Resolving Plant Phenology
@@
Figure 10: (Top) Query & Verification. Filtering top INQUIRE-Search outputs for "flowering" phenological stage. (Bottom) Phenological stages occur at significant times throughout the year. (left) density of observations from INQUIRE-Search corresponding to different stages. (right) distribution of observations from INQUIRE-Search in different phenological stages visualized with results from ANOVA/Tukey test.
Results and Analysis
@@
This case study shows that INQUIRE-Search can be used systematically to resolve detailed phenological information at regional scales for specific species. Some phenological stages, namely, flowering and seeding, were more successfully resolved than emergence and senescence. While this could be a limitation of the models ability to identify relevant observations to those stages (the difference between an emerging plant and a young plant is minor) it is also likely due to bias in the community science observations queried, as observations of flowers and seeds are likely more common as they can aid taxonomic identification [iNaturalist202xHighQualityObservations]. Resolving patterns of senescence will also require specific considerations to account for the fact that senescent plants can also be observed long after senescence, evidenced by the fact we identified observations of senescent milkweed throughout the year. Nonetheless, this sort of fine-scale phenological information would have traditionally required daily or weekly surveys to resolve at the level of a single site, let alone a region like Southern Quebec. Our results demonstrate that INQUIRE-Search has the potential to drastically increase our ability to extract phenological information from the wealth of rapidly growing community science observations.
3.5 C5: Whale Re-Identification
Results and Analysis